# Noise-based reward-modulated learning

### v1

@@ 4 Discussion
In this work, we introduced NRL as a novel noise-driven learning mechanism for neural networks that bridges RL principles with Hebbian-like updates, providing a gradient-free alternative to traditional approaches. By leveraging directional derivatives and stochastic noisy neurons to approximate gradients [44, 53, 52], our method produces local synaptic updates that are modulated by a global reinforcement signal. This noise-based rule offers an efficient approach for neural adaptation, particularly suited to real-time and resource-limited environments.
@@
NRL shares similarities with node perturbation methods [15, 20, 54, 52, 13], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation –which requires constant feedback or reward for learning– NRL is well-suited for delayed-reward setups, a common scenario in real-world applications. By leveraging eligibility traces [19], NRL can effectively assign credit to past neural states even when actions and outcomes are separated in time.
Furthermore, NRL shares conceptual similarities with forward gradient methods [2], which also estimate gradients using directional derivatives, stochastic perturbations, and forward-mode differentiation. Both frameworks leverage noise-driven updates, focusing on solving real-world tasks. However, while forward gradient methods rely solely on random perturbations to approximate gradients, NRL refines updates by incorporating reinforcement learning principles, using RPEs to directly modulate synaptic weights. Recent advances in forward gradient techniques, such as employing local auxiliary networks to produce informed gradient guesses [16], have demonstrated substantial improvements in alignment with exact gradients, narrowing the performance gap with backpropagation. Similarly, these variance-reduction strategies could inspire further refinements in our approach, enhancing the quality of noise-based updates and improving scalability to more complex architectures.
Another increasingly popular approach for gradient estimation during optimization is zeroth-order (ZO) optimization [28, 6]. Like NRL, common ZO optimization relies on perturbation-driven strategies, utilizing differences in function values resulting from input perturbations to approximate gradients. However, as with forward gradient methods, NRL distinguishes itself by incorporating elements, such as RPEs and eligibility traces, to learn in delayed reward scenarios. Unlike forward gradient methods, which rely on directional derivatives and access to the model’s structure, ZO methods estimate gradients purely from function evaluations, without requiring any knowledge of the network’s internal structure or differentiability. This makes ZO methods particularly well-suited for black-box optimization tasks, a flexibility that our approach also shares.
The results demonstrate the effectiveness of NRL, significantly outperforming the comparable RMHL approach [26, 32] in both instantaneous and delayed reward tasks. While RMHL methods aim to align neural dynamics with observed brain dynamics rather than optimizing learning, our learning rule is rigorously derived to perform gradient approximation. As a result, our learning rule introduces two key innovations. First, normalizing the noise, $\bar{\xi}^{l}_{t}=\ifrac{\xi^{l}_{t}}{||\xi^{l}_{t}||^{2}}$, allows for smaller noise magnitudes that minimally impact agent behavior while still contributing effectively to the learning process. Second, an additional term, ${p}_{t}=\left[\log\pi_{\tilde{\theta}}(a|S_{t})-\log\pi_{\theta}(a|S_{t})\right]$, captures the influence of noise on the network output, thereby enhancing credit assignment accuracy.
Our results further demonstrate that NRL effectively assigns credit even in deep networks with multiple hidden layers, a challenge for many biologically plausible algorithms. As network depth increases, NRL’s performance advantage over RMHL becomes more pronounced. In fact, RMHL struggles with credit assignment in networks with more than one hidden layer, as evidenced by its minimal performance improvements during training. Nevertheless, NRL does require more trials to converge in deeper networks, which is an expected consequence of its stochastic update mechanism [20].
Compared to traditional BP-based methods, NRL approximates gradients rather than relying on exact gradients [27, 51, 39, 38], offering a significant advantage in computationally constrained scenarios—such or when using non-differentiable networks like spiking neural networks [55, 50]. The local-update structure of NRL provides compatibility with neuromorphic hardware [42], where efficient, low-power computations are essential. Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key.
@@
The inherent noise in NRL is also advantageous for RL, as it naturally facilitates exploration-exploitation balance [22, 25]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
Our work further draws inspiration from neuroscience and underscores the importance of biologically plausible approaches [56, 30] for advancing learning algorithms. The reliance on RPEs as learning signals emulates the dopaminergic reward signals observed in the brain [41, 40], crucial for synaptic plasticity. The noise-based rule, coupled with eligibility traces, also aligns with biological processes that enable learning under uncertainty [14] and reflect the noisy nature of the brain [9, 45, 31]. Notice that the definition of RPEs used in this work is a simplification, relying on a straightforward running average for reward prediction. While this is sufficient to capture the core ideas of the proposed learning mechanism, more sophisticated approaches could further enhance reward estimation. One such extension is Temporal Difference (TD) learning [46, 48], which provides a more principled method for predicting expected future rewards and aligns more closely with biological dopaminergic activity [35]. A natural implementation would be an actor-critic architecture, where a critic network estimates the value function (e.g., expected future rewards), refining the reward prediction error (RPE).
One limitation of NRL is its dependence on multiple forward passes per synaptic update, which, though parallelizable, reduces biological plausibility. Future work could explore efficient ways to estimate noise impacts without explicitly needing separate clean and noisy passes. Additionally, while BP may offer faster convergence in unrestricted computational settings, recent research, including this study and others [7, 11], show promise in bridging this gap for specific networks and tasks.
Our findings suggest that noise-based learning strategies could drive advances in machine learning by enabling efficient, local, and energy-conscious algorithms, particularly suited for neuromorphic hardware and real-time applications. This research also offers valuable insights into how learning might occur in the brain, enriching our understanding of both artificial and biological intelligence.
@@ Appendix A Gradient approximation using directional derivatives
Consider a function $f(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
@@
The gradient $\nabla f$ is a vector indicating the direction of the steepest ascent of the function $f$, with components as partial derivatives of $f(\theta)$:
$\nabla f=\left[\frac{\partial f}{\partial\theta}\right]^{\top}=\left[\frac{\partial f}{\partial W_{1}},\ldots,\frac{\partial f}{\partial W_{L}}\right]^{\top}\,.$ (27)
While $\nabla f$ captures the rate of change of $f$ in the steepest direction, a directional derivative gives the rate of change in a specified direction. For a unit vector $v=\ifrac{\epsilon}{||\epsilon||}$, normalized via the Euclidean norm, we define the directional derivative as:
@@
The directional derivative of $f(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
$\nabla_{v}{f}(\theta)=\lim_{h\to 0}{\frac{f(\theta+hv)-f(\theta)}{h}}\,,$ (28)
@@
For a sufficiently small $h$ we approximate the directional derivative as
$\nabla_{{v}}{f}(\theta)=\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,.$ (29)
This directional derivative is essentially a projection of $\nabla f$ in the direction $v$, following the relation:
$\nabla_{v}{f}(\theta)=v\cdot\nabla{f}(\theta).$ (30)
@@
Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$, drawn from some probability distribution $p(\epsilon)$, and $n$ the number of dimensions in $\theta$. Exact gradients can be written in terms of directional derivatives using expectations
$\nabla f(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,.$ (31)
@@
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv\nabla f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv\right]\nabla f(\theta)^{\top}\,$
$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]\nabla f(\theta)\,$
$=n\frac{1}{n}I_{n}\nabla f(\theta)=\nabla f(\theta)\,$
as we assume $p(\epsilon)=\mathcal{N}(0,\sigma^{2}I_{n})$, which gives $\mathbb{E}[vv^{\top}]=\mathbb{E}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]=\frac{1}{n}I_{n}$. ∎
Now, consider gradient descent in the direction of the gradient $\nabla f(\theta)$ as
$\theta_{t+1}=\theta_{t}+\alpha\nabla f(\theta_{t}).$ (32)
@@
Let $\eta=\alpha n$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
$\theta_{t+1}=\theta_{t}+\alpha\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$ (33)
in the limit when $\sigma^{2}\to 0$ and $K\rightarrow\infty$.
@@
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[v\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
@@
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||^{2}}\left[f(\theta+\epsilon)-f(\theta)\,\right]^{\top}\right]\,$
We may approximate the expectation by sampling under some empirical distribution to obtain
$\nabla f(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$
In practice, sufficiently small $\sigma^{2}$ and limited noise samples approximate the gradient. For $K=1$, analogous to single-sample updates in stochastic gradient descent:
$\theta_{t+1}=\theta_{t}+\eta\frac{\epsilon}{||\epsilon||^{2}}\delta f$ (34)
with $\delta f=\left[f(\theta+\epsilon)-f(\theta)\right]^{\top}$.
@@ Appendix B Model hyperparameters and training details
[…]

### v2

@@ 4 Discussion
In this work, we propose NRL, a noise-driven synaptic plasticity rule that bridges principles of reinforcement learning and optimization with biologically-inspired Hebbian updates. By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal. It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
We employ a top-down methodology, starting by deriving our learning rule from the mathematical principles of gradient-based optimization. We then approximate the gradients via directional derivatives and incorporate bottom-up constraints inspired by biological systems. Specifically, the gradient-approximation process is implemented through stochastic neural activity (aligning with the inherent noise in the brain [12, 51, 39]), and reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [48, 47]), eligibility traces, and local synaptic plasticity. These elements enable learning under uncertainty [18]. This approach results in a learning rule that is both theoretically grounded and potentially realizable in neural circuits.
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. Additionally, it shows a significant improvement over RMHL [32, 40], a modern, well-known approach that employs noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible algorithms, indicating its potential for scaling. This is a critical difference with RMHL, which struggles to assign credit in multilayer networks evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
@@
NRL shares similarities with node perturbation methods [19, 24, 61, 58, 16], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation – which requires constant feedback or reward for learning – NRL is well-suited for delayed reward setups. By leveraging eligibility traces [23, 26], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable. Techniques such as forward gradient methods [2] and zeroth-order (ZO) optimization [35, 8] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients. Recent advances, such as employing local auxiliary networks for informed gradient guesses [20], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation. These variance-reduction strategies could similarly enhance our approach by improving the quality of the updates and scalability to more complex architectures. Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [62, 56].
@@
Additionally, the local-update structure of NRL provides compatibility with neuromorphic substrates [49], where efficient, low-power computations are essential. Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates exploration-exploitation balance [27, 30]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
Our work has limitations that suggest future research directions. First, its dependence on multiple forward passes per synaptic update reduces biological plausibility, even though these passes could be performed in parallel. Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes. Second, our modeling approach deliberately involves a high degree of abstraction: we use rate-based neurons instead of spiking models, do not distinguish between excitatory and inhibitory populations, and treat synaptic potentiation and depression as a single continuum. While this provides clear insight into the capabilities and limitations of our approach, more biologically detailed models could offer further insights. Third, our reward prediction error calculation is a simplification. While sufficient for demonstrating the learning mechanism, the simplicity of using a running average may limit its effectiveness in scenarios where navigating through transient low-reward regions is necessary to achieve higher long-term rewards. In such cases, the model may underestimate the value of actions that lead to delayed benefits, potentially resulting in suboptimal updates. More sophisticated approaches, such as temporal difference learning [52, 54], or learned prediction through a critic network, could provide a richer and more temporally sensitive estimate of expected rewards. These strategies, which align more closely with biological dopaminergic signaling and can be implemented within actor-critic architectures, represent promising directions for improving learning performance in future work. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [9, 14], including ours, indicate that this gap can be narrowed for specific networks and tasks.
Our findings highlight the potential of unifying optimization methods and neurobiological principles. By drawing inspiration from both domains, we have developed a learning rule that not only offers a biologically plausible alternative to backpropagation but also provides insights into how learning, particularly synaptic plasticity, might occur in biological neural systems. Conversely, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms, particularly suited for neuromorphic hardware and real-time applications. This synergistic approach promises to advance our understanding of intelligence in both biological and artificial systems [64, 38], driving progress in neuroscience and machine learning alike.
@@ Appendix A Gradient approximation using directional derivatives
Consider a function $g(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
@@
The gradient $\nabla g$ is a vector indicating the direction of the steepest ascent of the function $g$, with components as partial derivatives of $g(\theta)$:
$\nabla g=\left[\frac{\partial g}{\partial\theta}\right]^{\top}=\left[\frac{\partial g}{\partial W_{1}},\ldots,\frac{\partial g}{\partial W_{L}}\right]^{\top}\,.$ (22)
[…]
@@
The directional derivative of $g(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
$\nabla_{v}{g}(\theta)=\lim_{h\to 0}{\frac{g(\theta+hv)-g(\theta)}{h}}\,$, (23)
@@
For a sufficiently small $h$ we numerically measure the directional derivative as
[…]
This directional derivative can also be measured as a projection of $\nabla g$ in the direction $v$, following the relation:
$\nabla_{v}{g}(\theta)=v\cdot\nabla{g}(\theta).$ (25)
@@
Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ (probability distribution $p(\epsilon)$) where $n$ is the number of dimensions in $\theta$. Exact gradients can be written in terms of directional derivatives using expectations
$\nabla g(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,.$ (26)
@@
$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\nabla g(\theta)\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\right]\nabla g(\theta)\,$
[…]
$=n\frac{1}{n}I_{n}\nabla g(\theta)=\nabla g(\theta)\,$
[…]
Now, consider gradient descent in the direction of the gradient $\nabla g(\theta)$ as
$\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta_{t}).$ (27)
@@
Let $\eta=\frac{\alpha n}{hK}$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
[…]
in the limit when $h\to 0$ and $K\rightarrow\infty$.
@@
$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}g(\theta)^{\top}\right]\,$
@@
[…]
Now, substituting the expectation with a sampling under some empirical distribution,
[…]
Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$,
$\theta_{t+1}$ $=\theta_{t}+\alpha\nabla g(\theta)$
[…]
In practice, small perturbations, controlled by $h$, and limited numbers of noise samples are sufficient to approximate the gradient. For $K=1$, analogous to single-sample updates in stochastic gradient descent:
[…]
with $\rho=\left[g(\theta+h\epsilon)-g(\theta)\right]^{\top}$.
@@ Appendix B Model hyperparameters and training details
Acrobot $\eta$ 5e-3 5e-2 5e-2
$\sigma$ - 1e-3 1e-3
Cartpole $\eta$ 5e-3 5e-2 1e-2
$\sigma$ - 1e-3 1e-1
Reaching $\eta$ 1e-2 1e-2 1e-1
$\sigma$ - 1e-3 1e-1